A map has a scale of 3 cm : 7 km. If two cities are 8 cm apart on the map, what is the actual distance between the cities, to the nearest tenth of a kilometer?
Answer by stanbon(75887)

A map has a scale of 3 cm : 7 km. If two cities are 8 cm apart on the map, what is the actual distance between the cities, to the nearest tenth of a kilometer?
----
Use a proportion::
x/8 cm = 7km/3cm
----
x = 8(7/3) km
x = 56/3 km
x = 18.7 km (distance between the cities)
------------------
Cheers,
Stan H.
------------